# Deep Learning Based Segmentation of Blood Vessels from H&E Stained Oesophageal Adenocarcinoma Whole-Slide Images

### v1

@@ Abstract
Blood vessels (BVs) play a critical role in the Tumor Micro-Environment (TME), potentially influencing cancer progression and treatment response. However, manually quantifying BVs in Hematoxylin and Eosin (H&E) stained images is challenging and labor-intensive due to their heterogeneous appearances. We propose a novel approach of constructing guiding maps to improve the performance of state-of-the-art segmentation models for BV segmentation, the guiding maps encourage the models to learn representative features of BVs. This is particularly beneficial for computational pathology, where labeled training data is often limited and large models are prone to overfitting. We have quantitative and qualitative results to demonstrate the efficacy of our approach in improving segmentation accuracy. In future, we plan to validate this method to segment BVs across various tissue types and investigate the role of cellular structures in relation to BVs in the TME.
@@ 2.2 Guiding map algorithm
Fig.2 illustrates the steps involved in generating a guiding map, and Fig.3 shows examples of the resulting guiding maps.
@@ 3.3 Comparative experiments
To evaluate the effectiveness of our guiding map, we performed comparative experiments using the same architecture family (i.e.U-Net). We first trained the networks using RGB input, then we trained the same networks using 4-channel input, where guiding maps were concatenated to the RGB images. In our experiments, we trained U-Net models with different encoders: EfficientNet B0, B2, B6 [7] and EfficientNetV2-M [8] (we name these EfficientUnet). We incorporated Spatial and Channel Squeeze & Excitation (SCSE) blocks [9] into the decoder layers of the EfficientUnet models, these models were implemented using the Segmentation Models PyTorch library [10]. Basic U-Net[3], FCN[5] and TransUnet(R50-ViT-B_16)[11] were also trained on our dataset to establish benchmarks.
@@ 4 Results and Discussion
The test dataset contains both normal BVs and tumor BVs. Qualitative results, as shown in Fig.4, demonstrate that models trained with RGB + guiding map input achieved superior segmentation quality compared to those trained with RGB input alone. These models produced smoother segmentation masks, with fewer irregularities and a reduction in false positive predictions while being more sensitive to tumor BVs.
@@
A more detailed examination reveals that the use of guiding maps significantly increases the performance of EfficientUnet/B2 model in segmenting tumor BVs with an average increase of nearly 5% in DSC/IoU. This improvement was also observed across most models, except for EfficientUnet/B6 where the performance remains approximately the same, These results are reported in Table 2. For segmenting normal BVs, all models demonstrated improved performance when RGB + guiding map input was used, as shown in Table 3.
These findings suggest that the integration of our guiding maps enhances the ability of deep learning models to capture representative features of both normal and tumor BVs, especially in models with higher number of parameters, this further highlights the utility of the guiding maps in reducing overfitting, particularly when training data is limited and/or contains significant variability.
@@ 5 Conclusion and future work
We introduced a novel image processing algorithm to enhance deep learning-based segmentation of blood vessels (BVs) in H&E-stained oesophageal adenocarcinoma whole-slide images. By generating high-contrast guiding maps and concatenating them to RGB images, our method improves segmentation performance while requiring minimal changes to network architectures. Experimental results showed significant quantitative and qualitative improvements, especially in larger models, and when training data is limited, this highlights the value of guiding maps in enhancing the ability of models to capture representative features of BVs.
@@ 7 Acknowledgments
SA is supported by the NIHR Imperial Biomedical Research Centre, Cancer Research UK, and the Rosetrees Trust (ref: PGL23/100123). JL is funded by the UK Engineering and Physical Sciences Research Council (EPSRC). The authors acknowledge the use of the Batch Compute System in the Department of Computer Science at the University of Warwick, and associated support services, in the completion of this work.

### v2

@@ Abstract
Blood vessels (BVs) play a critical role in the Tumor Micro-Environment (TME), potentially influencing cancer progression and treatment response. However, manually quantifying BVs in Hematoxylin and Eosin (H&E) stained images is challenging and labor-intensive due to their heterogeneous appearances. We propose a novel approach of constructing guiding maps to improve the performance of state-of-the-art segmentation models for BV segmentation, the guiding maps encourage the models to learn representative features of BVs. This is particularly beneficial for computational pathology, where labeled training data is often limited and large models are prone to overfitting. The quantitative and qualitative results demonstrate the efficacy of our approach in improving segmentation accuracy. In future, we plan to validate this method to segment BVs across various tissue types and investigate the role of cellular structures in relation to BVs in the TME.
@@ 2.2 Guiding map algorithm
Fig. 2 illustrates the steps involved in generating a guiding map. Fig. 3 shows examples of the resulting guiding maps.
@@ 3.3 Comparative experiments
To evaluate the effectiveness of our guiding map, we performed comparative experiments using the same architecture family (i.e. U-Net). We first trained the networks using RGB input, then we trained the same networks using 4-channel input, where guiding maps were concatenated to the RGB images. In our experiments, we trained U-Net with different encoders: EfficientNet B0, B2, B6 [7] and EfficientNetV2-M [8] (we name these EfficientUnet). We incorporated Spatial and Channel Squeeze & Excitation (SCSE) blocks [9] into the decoder layers of the EfficientUnet models, these models were implemented using the PyTorch library [10]. Basic U-Net [3], FCN [5] and TransUnet (R50-ViT-B_16) [11] were also trained on our dataset to establish benchmarks.
@@ 4 Results and Discussion
The test dataset contains both normal BVs and tumor BVs. Qualitative results, are shown in Fig.4 which demonstrate that the models trained with RGB + guiding map achieve superior segmentation quality compared to those trained with RGB input only. These models produced smoother segmentation masks, with fewer irregularities and a reduction in false positive predictions while being more sensitive to tumor BVs.
@@
Further detailed examination reveals that the use of guiding maps significantly increases the performance of EfficientUnet/B2 model in segmenting tumor BVs with an average increase of nearly 5% in DSC/IoU. This improvement was also observed across other models, except for EfficientUnet/B6 where the performance remains approximately the same (see Table 2). For segmenting normal BVs, all models demonstrated improved performance when RGB + guiding map input was used, as shown in Table 3.
These findings suggest that the integration of the proposed guiding map enhances the ability of deep learning models to capture representative features of both normal and tumor BVs, especially in models with higher number of parameters. This further highlights the utility of the guiding maps in reducing overfitting, particularly when training data is limited and/or contains significant variability.
@@ 5 Conclusion and future work
We introduced a novel image processing algorithm to enhance deep learning-based segmentation of blood vessels (BVs) in H&E-stained oesophageal adenocarcinoma whole-slide images. By generating high-contrast guiding maps and concatenating them to RGB images, our method improves segmentation performance while requiring minimal changes to network architectures. The experimental results show significant improvements, especially in larger models, and when training data is limited. This highlights the significance of guiding maps in improving the ability of models to segment BVs.
@@ 7 Acknowledgments
SA is supported by the NIHR Imperial Biomedical Research Centre, Cancer Research UK, and the Rosetrees Trust (ref: PGL23/100123). JL is supported by the UK Engineering and Physical Sciences Research Council (EPSRC). SEAR report financial support provided by the MRC (MR/X011585/1) and the BigPicture project, funded by the European Commission.